I am having difficulty understanding this: find an equation of the line passing through (-2,4) and (6,-9). What is the first step to solving this problem?
Answer by htmentor(1343)

I am having difficulty understanding this: find an equation of the line passing through (-2,4) and (6,-9). What is the first step to solving this problem?
=============================
The slope of the line, m = (y2-y1)/(x2-x1) = (-9 - 4)/(6 - -2) = -13/8
The standard form for a line is y = mx + b
Use one of the (x,y) pairs to solve for b:
4 = (-13/8)(-2) + b
b = 32/8 - 26/8 = 6/8 = 3/4
So the equation for the line is y = (-13/8)x + 3/4